If the position vector of A and B are 3 −→a − 7 −→b − 7 −→c and 5 −→a + 4 −→b + 3−→c , nd −→AB and determine its magnitude and direction cosines.


1
Expert's answer
2022-02-17T15:41:55-0500
"\overrightarrow{AB}=5\vec a+4\vec b+3\vec c-(3\vec a-7\vec b-7\vec c)"

"=2\vec a+11\vec b+10\vec c"

"|\overrightarrow{AB}|=\sqrt{2^2+11^2+10^2}=15"


"\cos \alpha=\dfrac{2}{15}"

"\cos \beta=\dfrac{11}{15}"

"\cos \gamma=\dfrac{10}{15}"